A uniform chain of length lvand mass m lies on a smooth table .a very small lart of it this chain hangs from table .it begins to fall freely under the weight of hanging chainvfind the velocity of chain when the length of hanging part becomes y

Grade:12

1 Answers

Manas Shukla
7 years ago
Length of the chain = L
Given the chain is of uniform density and shape.
Let mass of chain = M
Hanging part of the chain = y
depth of centre of mass of the chain of hanging part = y/2
Since the chain is rigid it will move with the same speed.
Mass of the hanging part of the chain = My/L
Lets balance Energy when hanging part becomes y.
\frac{1}{2} \times M\times v^{2} = \frac{My}{L}\times g\times \frac{y}{2}
v = y\sqrt{\frac{g}{L}}